6.80 mL of a 1.30 M solution of lead(II) nitrate was combined with 3.20 mL of a 5.0 M
solution of ammonium sulfide.
1. Write the chemical equation for this reaction and balance it.
2. What is the total concentration of all soluble ions after these two solutions are mixed
and allowed to react completely?
3. A solid product precipitated out of this reaction, and was collected. If 2.01 g of this
product were collected, what is the percent yield?
Answer #1

1) The reaction will be

Here, PbS will precipitate out.

2)

Total Volume = 6.80 mL + 3.20 mL = 10.00 mL

milli moles of Pb(NO3)2 = Volume * Molarity = 6.80 mL * 1.30 M = 8.84 mmmol

milli moles of (NH4)2S = Volume * Molarity = 3.20 mL * 5.0 M = 16.0 mmmol

Concentration of NO3-

1 mmol of Pb(NO3)2 will give 2 mmol of NO3- which will remain soluble.

Thus, mmol of NO3- = 2 * 8.84 mmol = 17.68 mmol

Concentration of NO3- = Moles/Volume = 17.68 mmol / 10.00 mL = 1.768 M

Concentration of NH4+

1 mmol of (NH4)2Swill give 2 mmol of NH4+ which will remain soluble.

Thus, mmol of NH4+ = 2 * 16.00 mmol = 32.00 mmol

Concentration of NH4+ = Moles/Volume = 32.00 mmol/10.00 mL = 3.200 M

Concentration of Pb2+

1 mmol of Pb(NO3)2 will give 1 mmol of Pb2+

Thus, mmol of Pb2+= 8.84 mmol

But all of it will react with 8.84 mmol S2- to form precipitate and none of it will remain.

Concentration of Pb2+ =0 M

Concentration of S2-

1 mmol of (NH4)2Swill give 1 mmol of S2- which will remain soluble.

Thus, mmol of S2- = 16.00 mmol

But 8.84 mmol will react with Pb2+.

Moles remaining = 16.00 mmol - 8.84 mmol = 7.16 mmol

Concentration of S2-  = Moles/Volume = 7.16 mmol/10.00 mL = 0.716 M

Total concentration of all soluble ions = 1.768 M + 3.200 M + 0.716 M = 5.684 M = 5.68 M (3 sig figure)

c)

Theoretical moles of PbS produced = 8.84 mmol = 8.84 * 10-3 mol

Molar Mass of PbS = 239.3 g/mol

Theoretical Mass = (8.84 * 10-3 mol) * (239.3 g/mol) = 2.115 g

Actual Mass = 2.01 g

% Yield = Actual Mass / Theoretical Mass * 100 = 2.01 g / 2.115 g * 100= 95.03%